OpenStudy (anonymous):

(2x+3) / 4 - (3x-5) / 3 = 2 solve for x.

OpenStudy (psymon):

First thing I would do is get rid of those denominators. You can get rid of something in the denominator my multiplying everything by that same number. So for example, if I wanted to get rid of that 4 in your first denominator, I would simply multiply everything by 4: \[\frac{ 4(2x+3) }{ 4 }-\frac{ 4(3x-5) }{ 3 }=(2)(4)\]Note how I multiplied every term by 4. You must make sure everything gets multipliedif you do this. Now those 4's on that left fraction will cancel out and I will have this: \[(2x+3)-\frac{ 4(3x-5) }{ 3 }=8\] Also note how I keep thoseparenthesis. It's a very good habit to have, it'll help make sure you do not get the wrong signs or forget to do some sort of operation. Now I have that 3 left in the 2nd denominator. I can eliminate that in the same method I did with the 4, by multiplyong every by 3 like so: \[3(2x+3) - \frac{ (3)(4)(3x-5) }{ 3 }=(8)(3)\]Again, I was careful to multiply everything by 3. Also notice how I just leave a bunch of things grouped together, but not actually multiplied out. This is also sometimes a good strategy. It lets you visualize what you're doing better in case you miss something or need to backtrack. Now, I can cancel out that 3 in the second denominator, which will leave me with this. \[3(2x+3)-4(3x-5)=24\] That was probably the trickiest part of this problem. I'll leave it to you to see if you can finish the rest from here : )

OpenStudy (anonymous):

but we have to do it this way for the test: can you show me where i am going wrong?\[\frac{ 2x+3 }{ 4} - \frac{ 3x-5 }{ 3 } = 2\]

OpenStudy (anonymous):

then do this: \[\frac{ 3(2x+4) }{ 12 } - \frac{ 4(3x-5) }{ 12 } = 2\]

OpenStudy (anonymous):

then: \[\frac{ 6x+12 - 12x - 2x +20 }{ 12 }\]

OpenStudy (anonymous):

where am i going wrong??

OpenStudy (anonymous):

@daftkillz - where did u get the "-2x" term in this equation ??

OpenStudy (anonymous):

sorry delete that

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

pretend its not there

OpenStudy (anonymous):

its still wrong -_-

OpenStudy (anonymous):

also it should be 3(2x+3)

OpenStudy (campbell_st):

your error is in distributing 3(2x + 3) you have 6x + 12..... which isn't correct

OpenStudy (anonymous):

so it should be: 6x+9 - 12x +20 = 2

OpenStudy (anonymous):

because thats still wrong..

OpenStudy (anonymous):

RHS should be 12*2=24 right ?

OpenStudy (anonymous):

wait what?

OpenStudy (campbell_st):

well if you change the 2 to 24... multiply every term by 12 you can solve your equation

OpenStudy (anonymous):

wow made a stupid as mistake ok i get it now

OpenStudy (anonymous):

i forgot 2 multiply the rhs by 12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Amy is twice as old as lucy. Three years from now the sum of their ages will be 42. how old is amy?

OpenStudy (anonymous):

how do you do this?

OpenStudy (anonymous):

can u share the equations u wrote for this problem ?

OpenStudy (anonymous):

i did this: Let lucy= L and amy= A 3L + 3 = 42 3L = 39 2L = 26 therefore amy = 26 but the answer is 24

OpenStudy (anonymous):

???

OpenStudy (anonymous):

do it one by one... also you forgot to add three years to lucy's or amy's age :P

OpenStudy (anonymous):

thts all

OpenStudy (anonymous):

ohhh ok thanks :D